Solve for x:
x^2.5 = 15.758

Answer by Boreal(15235)

x^2.5=15.758
2.5 log x=log 15.758
log x=log 15.758/2.5
log x=1.197/2.5. Without rounding, then take 10 and raise it to the power of the quotient. That removes the log and gets x.
x=3.013 rounding at the end.
3.013^2.5=15.75788.